OpenStudy (anonymous):

Please help, 2 equations

OpenStudy (anonymous):

1. Solve for b, then find the value of b when a = 3 4a = 2b - 7 A. -9/2 B. 5/2 C. 12/2 D. 17 2. Solve for r d = rt A. r = dt B. r = t/d C. d/t D. r = d - t

OpenStudy (anonymous):

So how would you isolate b in one side?

OpenStudy (anonymous):